A cricketer throws a 147g ball by exerting a force of 7.5N for 0.84s. If the launch angle is 39° from horizontal, calculate the range of the ball.
1
Expert's answer
2020-01-08T09:48:39-0500

Find initial velocity


"v_0=\frac{F}{m}t_0"

It's known problem - throwing the body at an angle to the horizon


"x=v_0cos(\alpha)t""y=v_0sin(\alpha)t-\frac{gt^2}{2}"

From this equation


"x=\frac{2v_0^2sin(\alpha)cos(\alpha)}{g}"

Calculating


"x=\frac{2\cdot (\frac{7.5\cdot 0.84}{0.147})^2 \cdot sin(39)\cdot cos(39)}{10}=183 (m)"